OpenStudy (anonymous):

(-2a^2)^3 = ? i believe its -8a^6 but not certain so just checking my answer

OpenStudy (kainui):

Correctomundo

OpenStudy (anonymous):

2^3=8

OpenStudy (anonymous):

(a^2)^3=4^6

OpenStudy (anonymous):

*a^6 not 4^6

OpenStudy (kainui):

Just tell him he's right for god"s sake @L-Lawliet-L ! READ!

OpenStudy (anonymous):

He's right.

OpenStudy (austinl):

For example\[(x)^3(x)^4 = (xxx)(xxxx) = x^7\] (-2)(-2)(-2) = -8 You add the exponents, which is what you did. So correct. :D

OpenStudy (luigi0210):

Well he gave an explanation, which you're suppose to be doing anyways.

OpenStudy (austinl):

Not if it is just verification of their work. If they are correct, what purpose would it serve to further explain a question that they already have correct.